# Strategic Tradeoffs Between Humans and AI in Multi-Agent Bargaining

### v3

@@
Abstract
As large language models (LLMs) are increasingly embedded in collaborative human activities such as business negotiations and group coordination, it becomes critical to evaluate both the performance gains they can achieve and how they interact in dynamic, multi-agent environments. Unlike traditional statistical agents such as Bayesian models, which may excel under well-specified conditions, large language models (LLMs) can generalize across diverse, real-world scenarios, raising new questions about how their strategies and behaviors compare to those of humans and other agent types. In this work, we compare outcomes and behavioral dynamics across humans (N = 216), LLMs (GPT-4o, Gemini 1.5 Pro), and Bayesian agents in a dynamic negotiation setting under identical conditions. Bayesian agents extract the highest surplus through aggressive optimization, at the cost of frequent trade rejections. Humans and LLMs achieve similar overall surplus, but through distinct behaviors: LLMs favor conservative, concessionary trades with few rejections, while humans employ more strategic, risk-taking, and fairness-oriented behaviors. Thus, we find that performance parity—a common benchmark in agent evaluation—can conceal fundamental differences in process and alignment, which are critical for practical deployment in real-world coordination tasks. By establishing foundational behavioral baselines under matched conditions, this work provides a baseline for future studies in more applied, variable-rich environments.
1 Introduction
Historically, bespoke statistical models have excelled in structured economic exchanges with well-defined action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). More recently, the generalizability and flexibility of large language models (LLMs) have enabled delegation in unstructured contexts: negotiating supply chain contracts (Van Hoek et al., 2022), coordinating international trading (Corvin, 2024), and facilitating online commerce transactions (PYMNTS, 2025; Gaarlandt et al., 2025). Unlike more structured models, LLMs can generalize under uncertainty and partial information, reducing the need for task-specific customization. As AI agents grow more sophisticated and are entrusted with higher-stakes decisions, understanding how different types of agents behave in dynamic, multi-agent settings is essential. In particular, navigating the trade-offs between agent types—not only in terms of negotiated outcomes, but also in their alignment with human strategies and norms—is critical for responsible delegation.
@@ 1 Introduction
A barrier to understanding delegation tradeoffs has been the difficulty of evaluating disparate agent types in comparable interactive settings. Static QA-style benchmarks fail to capture the dynamic, multi-agent interactions that characterize real-world negotiation. Indeed, recent work has called for more realistic evaluation environments that reflect the social and strategic complexities faced by deployed AI agents (Raman et al., 2024; Goktas et al., 2025).
We address this challenge by introducing a novel multi-player bargaining game that enables direct comparison of different agent types (e.g., humans, LLMs, and statistical models) under identical conditions. In our game, players receive endowments of colored chips with private valuations and negotiate trades to maximize their individual surplus. Bargaining is an ideal testbed for dynamic, multi-agent evaluations, as it inherently exposes tensions between social norms, individual preferences, and strategic reasoning, reflecting the complexities of real-world environments. Moreover, from a theoretical perspective, bargaining under partial information admits no Pareto efficient, incentive-compatible solutions (Myerson and Satterthwaite, 1983a), forcing agents to navigate trade-offs between individual gain and mutually beneficial cooperation.
We conduct an empirical study of our game across three types of agents: (i) human participants (N = 216), (ii) LLM-based agents, and (iii) specialized Bayesian agents, all facing identical randomized endowments. The Bayesian agents achieved the highest surplus, proposing aggressive bargains at the cost of frequent rejections. LLMs matched human surplus in certain game conditions but exhibited fundamentally different behaviors: where humans made fairness-driven concessions, LLMs favored conservative, low-variance trades with high acceptance rates.
More broadly, these findings expose a limitation of outcome-based evaluation: aggregate surplus can mask important differences in behavior and alignment. As AI systems increasingly mediate real-world negotiations, it becomes essential to assess not just what agents achieve, but how they navigate trade-offs, adjust to strategic dynamics, and engage with social norms. The primary contributions of this work are threefold:
@@
  • A dynamic multi-agent environment for comparative evaluation. We design and implement a bargaining game that enables direct, identical-condition comparisons of negotiation strategies across different agent types, with game variations of increasing complexity. We accompany the platform with a linear programming formulation of the Pareto-optimal surplus, providing a theoretical benchmark for performance.
  • An empirical characterization of human, LLM, and Bayesian negotiation strategies. Through a study with 216 human participants (N=216), we reveal distinct behavioral profiles: Bayesian agents adopt aggressive, surplus-maximizing strategies with high rejection rates; humans exhibit strategic concessions and fairness norms; LLMs act as conservative agents that yield high rates of offer acceptance.
@@
  • A framework to demonstrate how performance parity can conceal strategic misalignment. Our experimental benchmarks show that aggregate outcomes (i.e., total surplus) are insufficient for evaluation, as they mask fundamental differences in agent behavior. This finding highlights the need to move beyond performance-based metrics to assess process and alignment for the responsible deployment of agents in real-world coordination tasks.
2 Related work
@@
Statistical agents have excelled in well-defined tasks with explicit objectives, preferences, and constrained action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). Recent advances in LLMs have broadened the scope of delegation to include tasks requiring broader social reasoning, such as preference elicitation, strategic pricing, natural-language negotiation, and alliance formation in multi-agent games like Diplomacy (Fish et al., 2024; Soumalias et al., 2025; Tessler et al., 2024; Vaccaro et al., 2025; Lubars and Tan, 2019; Bakhtin et al., 2022; Kramár et al., 2022; Deng et al., 2024). LLMs have exhibited emergent capabilities in economic computation and human alignment, replicating behaviors in bargaining, auctions, and other economic games (Horton, 2023; Aher et al., 2022; Argyle et al., 2022), and reflecting human subgroup-specific patterns when conditioned on demographic traits (Manning et al., 2024; Xie et al., 2024; Zhu et al., 2024; Park et al., 2023; Qian et al., 2025).
@@
However, appropriate deployment and delegation depend not only on system capabilities (Jahani et al., 2025), but also on user expectations, skills, and alignment with social norms; users may suboptimally delegate when system behavior violates expectations (Mozannar et al., 2023; Agarwal et al., 2023; Mullainathan and Obermeyer, 2021; Palminteri et al., 2024; Kapania et al., 2022; Qian and Wexler, 2024). Addressing alignment gaps requires evaluating both capabilities and behaviors in dynamic, social contexts—a need recognized by recent calls for rigorous methods to steer and assess agent behavior in strategic multi-agent settings (Raman et al., 2024; Goktas et al., 2025). To concretely examine these challenges, we study how humans, LLM-based agents, and Bayesian models behave in a dynamic bargaining task, offering a comparative lens on alignment in strategic social interactions.
3 The bargaining game
3.1 Motivation from real-world negotiations
Negotiation remains one of humanity’s most fundamental coordination tasks Nash et al. (1950); Rubinstein (1982). Bilateral and multi-party bargaining appear in contexts such as three-team NBA tradesPelechrinis et al. (2023), B2B inventory swaps, and corporate barter networks Binmore (2007). These environments require participants to negotiate over bundles of heterogeneous assets with private valuations, making offers such as “X of A for Y of B (plus cash),” under real-world feasibility and time constraints.
Such bargaining is both strategically and communicatively demanding Farrell and Rabin (1996); Camerer (2003), motivating efforts to automate or augment the process. With advances in large language models (LLMs), it is now plausible to automate not just basic proposals, but also reasoning, persuasion, and adaptation in natural language. However, the strategic differences across LLM and human negotiation behavior remain underexplored; LLMs may diverge from human norms in ways that shape outcomes, affecting agreement rates, surplus allocation, and perceptions of fairness.
Studying these trade-offs in applied settings invites confounds; for example, if we were to evaluate the exchange of real commodities like secondhand goods in an online marketplace, participants’ heterogeneous preferences would confound direct comparison across agent types. To address this, we construct an abstract bargaining game that captures the essential structure of bargaining markets: fixed assets, a shared numeraire, private valuations, and an offer–response protocol Roth and Kagel (1995). The tensions embedded in our game—private information, strategic risk, and the trade-off between individual gain and mutual benefit—are fundamental to a wide range of real-world coordination tasks, from collective bargaining for labor contracts to negotiating international climate treaties Schelling (1980); Myerson and Satterthwaite (1983b). Therefore, our findings on agent behavior provide a foundational and generalizable diagnostic for understanding how these agents will likely act in more complex, unstructured interactions.
3.2 A stylized game
We introduce a multi-player bargaining game which captures strategic trade-offs, social signaling, and coordination challenges, influenced by real-world negotiations. Figure 1 presents an overview of the game, in which three players aim to maximize individual surplus over nine turns. Each begins with ten chips of each color. Green chips act as a shared numeraire valued at \$0.50 by all; other chip colors (red, blue, purple) have private values, drawn uniformly between \$0.10 and \$1.00. Game complexity is varied by incrementally adding chip types: red, then blue, then purple, with green always present. [^1] We show in Appendix A.2 that decision complexity grows with the number of chip colors.
Figure 1: This provides a broad overview of the bargaining game. Left: Game setup with initial chip counts and valuations. Center: Simplified version of gameplay (offer proposals and acceptances). Right: The types of agents empirically evaluated and simulated in this study.
@@ 3.2 A stylized game
In a fixed, random order, players propose trades by offering chip(s) of one color in exchange for another. The two non-proposers simultaneously decide whether to accept or decline the proposal; if both accept, one is chosen at random to clear the transaction. Players can only propose and accept offers that they have sufficient inventory to fulfill. All players observe transactions and chip holdings, but not others’ chip valuations. At the end of the game, each player receives a payout based on the surplus generated from their final chip holdings.
4 A pareto optimal upper bound
Before exploring empirical outcomes, we provide a theoretical upper bound analysis. Although our bargaining game lacks a tractable exact game-theoretic solution, we can compute a Pareto-optimal allocation via linear programming to benchmark agent performance in our empirical results.
4.1 Notation
Consider a game M = (I, G, v, a) with a set of agents I and chips G. Each agent $i$ has a personal valuation $v_{ig}$ for each good $g$. These valuations are static, independent across agents, and complete for each agent-good pair. All agents have an initial allocation $a^{0}_{ig}$ of each chip, which indicates the amount of chip $g$ allocated to agent $i$. Given these allocations, an agent’s welfare $w_{i}$ is the sum of their valuations weighted by their allocations: $w_{i}=\sum_{g\in G}v_{ig}a_{ig}$. The total welfare w is thus the sum of all individual welfare: $\textbf{w}=\sum_{i\in I}w_{i}=\sum_{i\in I}\sum_{g\in G}v_{ig}a_{ig}$. We will refer to the difference between initial and post-trade welfare as surplus gain in the following analysis.
4.2 Efficient allocations
In our game, agents trade chips to improve their welfare. Given their initial endowments $a^{0}_{ig}$, mutually beneficial trades may exist that increase total welfare. We provide an upper bound to the obtainable total welfare, assuming the Pareto condition that no player ends with worse utility than they began.
@@ 4.2 Efficient allocations
(Optimal Allocation of Chips). An optimal allocation of chips is any set of allocations $\mathbf{A^{*}}=\{a^{*}_{ig}\ |\ i\in I,g\in G\}$ that solves:
@@
Intuitively, in a centralized setting with full information, $\mathbf{A^{*}}$ represents a Pareto-efficient allocation that maximizes aggregate utility without making any agent worse off. While this optimal allocation $\mathbf{A^{*}}$ may not be unique, the maximum total welfare $\mathbf{w^{*}}$ achieved at the optimum is unique. This optimization problem can be solved efficiently using interior point methods. Its computational complexity is bounded by $O\bigl((|I||G|)^{3.5}\bigr)$. Note that this upper bound is unlikely to be achieved in practice, as no mechanism for decentralized bargaining with private information can guarantee both Pareto efficiency and incentive compatibility Myerson and Satterthwaite (1983a). [^2] A proof that players of the bargaining game lack a dominant strategy is provided in AppendixA.1.
5 Experimental setup
In our design, human participants first completed games in groups of three. Each group was randomly assigned to either the 2-chip, 3-chip, or 4-chip game variation, allowing us to evaluate performance and strategic trade-offs under increasing task complexity. We then replicated each game (identical initial chip endowments) with LLM- and Bayesian- agent groups, enabling direct comparison across the groups. [^3] For reproducibility, details on the human game interface are provided in Appendix B.1. Full LLM prompts are provided in Appendix G, and the Bayesian agent algorithms are detailed in Appendix A.3.
5.1 Human subjects

### v4

@@
Abstract.
Markets increasingly accommodate large language models (LLMs) as autonomous decision-making agents. As this transition occurs, it becomes critical to evaluate how these agents behave relative to their human and task-specific statistical predecessors. In this work, we present results from an empirical study comparing humans (N=216), multiple frontier LLMs, and customized Bayesian agents in dynamic multi-player bargaining games under identical conditions. Bayesian agents extract the highest surplus with aggressive trade proposals that are frequently rejected. Humans and LLMs achieve comparable aggregate surplus within their groups, but exhibit different trading strategies. LLMs favor conservative, concessionary proposals that are usually accepted by other LLMs, while humans propose trades that are consistent with fairness norms but are more likely to be rejected. These findings highlight that performance parity—a common benchmark in agent evaluation—can mask substantive procedural differences in how LLMs behave in complex multi-agent interactions.
multi-agent bargaining; negotiation; large language models; autonomous agents; strategic behavior; experimental study; Bayesian agents; fairness norms
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789078isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Multi-agent systemsccs: Theory of computation Algorithmic game theory and mechanism designccs: Computing methodologies Artificial intelligence
1. Introduction
For actions with real social or economic consequences, humans have traditionally been the decision-makers, while statistical models have handled narrowly-scoped, well-defined tasks. As large language models (LLMs) have advanced, they have begun to upend this paradigm (Eloundou et al., 2024; Shahidi et al., 2025), exhibiting strong zero- and few-shot performance across many domains (Brown et al., 2020). For example, call-center inquiries were once handled solely by human agents, then fronted by scripts and decision trees, and are now increasingly triaged by LLMs before reaching humans (Brynjolfsson et al., 2025). Similar shifts are emerging across other interpersonal tasks. LLMs are used to negotiate contracts (Van Hoek et al., 2022), coordinate trade (Corvin, 2024), and facilitate e-commerce (PYMNTS, 2025; Gaarlandt et al., 2025). This proliferation continues even though LLMs can fail to generalize in unexpected ways (Vafa et al., 2024). As they take on higher-stakes decisions, understanding their behavior relative to humans and previous generations of statistical models becomes essential—not only in terms of outcomes, but also in terms of alignment with procedural strategies and norms.
@@ 1. Introduction
Evaluating these tradeoffs requires assessing different types of agents in comparable yet sufficiently complex settings. This is difficult in practice. Across real-world applications, populations, incentives, and interfaces vary, making it hard to attribute observed differences to the agent rather than the setting. Static QA-style benchmarks—still the dominant LLM evaluation paradigm (Hendrycks et al., 2021)—fail to capture the dynamic, multi-agent settings that characterize many real-world interactions. Indeed, recent work has called for more realistic evaluation environments that reflect the social and strategic complexities of the real-world (Raman et al., 2024; Goktas et al., 2025).
We address this challenge by introducing an incentivized, multi-player bargaining game. The game is designed to enable controlled comparisons among LLMs, humans, and bespoke statistical agents under identical conditions. In our setting, players are endowed with colored chips associated with fixed, private valuations, and they negotiate trades to maximize individual surplus, defined as the value of their final chip bundle minus the value of their initial endowment. Crucially, preferences are fully induced (Smith, 1976): endowments and valuations are randomly drawn from a known distribution. This design ensures that differences in outcomes or behavior are attributable to differences in the agents themselves, rather than differences in underlying preferences.
This game has several features that make it well-suited for comparing agent types. First, its specific formulation is novel, providing a meaningful out-of-distribution domain for LLM agents. Second, despite the controlled structure of the game, bargaining under partial information admits no Pareto-efficient, incentive-compatible solution (Myerson and Satterthwaite, 1983), forcing agents to navigate genuine tradeoffs between self-interest and cooperation. Third, we can vary strategic complexity by adjusting the number of chip types in players’ endowments. Finally, while the game is not itself incentive-compatible, it does admit a computable Pareto-optimal upper bound, which can benchmark performance.
We conduct an empirical study of how different agent types play this game under identical conditions. First, we run an incentivized online study with human participants (N=216N=216). [^1] Humans are paid in USD per their gains from trade beyond their initial endowment. We then replicate the same randomized initial conditions in simulation with (i) LLM-based agents (GPT-4o and Gemini 1.5 Pro) and (ii) specialized Bayesian agents, enabling controlled comparisons across the three agent types. This setup enables an examination of two central research questions:
  1. RQ1:
    Performance: How does the performance of LLM agents compare to bespoke statistical agents and humans?
  2. RQ2:
    Process: How do the trading strategies and resulting group dynamics of LLM agents compare to those of other agents?
Figure 1. Overview of the bargaining game. Left: Game setup showing chip endowments and valuation structure. Center: Gameplay mechanics for proposing and accepting offers. Right: The three agent types evaluated in this study.
Three-panel overview of a three-player bargaining game. Left: players receive chips of 2–4 colors; chip counts are public; green has a shared value, and other colors have private random values. Center: players propose in a fixed cycle; the two others accept or reject, yielding no deal, a bilateral deal, or a random tie-break if both accept. Right: agents include humans, LLMs (Gemini 1.5 Pro, GPT-4o), and heuristic baselines (myopic, Bayesian).
We find that the game proves nontrivial: across all games and complexity levels, no agent type—human or artificial—achieves more than 80% of the Pareto bound on average. The agents’ performance varied substantially, and the processes by which they traded were remarkably different. Bayesian agents achieved the highest average surplus (up to 80% of the Pareto bound) by proposing highly self-interested trades that were often rejected. LLM agents were capable of achieving similar surplus (62%) to their human counterparts (59%). [^2] These statistics are for the medium-complexity game variation (3-chip). See Section 3.
Despite comparable outcomes, strategic behaviors between humans and GPT-4o agents diverged substantially. Humans offered trades that appeared more “fair,” while LLMs proposed more conservative trades that gave other LLMs within their group substantial surplus and were accepted at higher rates. Taken together, these results highlight the limitations of outcome-driven evaluation, as the similar surplus gains by humans and LLMs masked consequential differences in strategic behavior and procedural alignment.
In summary, this work makes three core contributions:
  1. 1.
@@ 2. RQ2:
    A dynamic multi-agent environment for comparative evaluation. We design and implement an open-source, incentivized bargaining game that enables ceteris paribus comparisons of behavior across agent types, with game variations of increasing complexity. The game is extensible to agent types not studied here (e.g., other LLMs). We also provide a linear programming formulation of the Pareto-optimal surplus as a theoretical benchmark for performance.
  2. 2.
    An empirical sample of human, LLM, and Bayesian trading strategies. We characterize how each agent type negotiates within a bargaining scenario under identical conditions: Bayesian agents aggressively maximize surplus and frequently reject offers; humans make strategic concessions motivated by fairness norms; LLMs negotiate conservatively and accept most offers.
  3. 3.
@@
    Evidence that outcome similarity can mask meaningful differences in behavior. We show that aggregate outcomes (e.g., total surplus) can conceal substantial differences in strategies, interaction patterns, and normative alignment. This finding highlights the need to move beyond performance-based metrics for the responsible deployment of LLMs in high-stakes real-world settings.
2. Related Work
@@
Recent advances in LLMs’ capabilities have dramatically broadened the scope of tasks that they can perform in scenarios involving collective reasoning, including understanding of social context, preference elicitation, strategic pricing, and natural-language negotiation (Fish et al., 2024; Soumalias et al., 2025; Tessler et al., 2024; Vaccaro et al., 2025; Lubars and Tan, 2019; Bakhtin et al., 2022; Kramár et al., 2022; Deng et al., 2024; Allouah et al., 2025). LLMs have also demonstrated the capacity to perform economic tasks, replicating and predicting behaviors in bargaining, auctions, and other games (Aher et al., 2022; Argyle et al., 2022; Horton, 2023; Manning et al., 2024; Manning and Horton, 2025; Xie et al., 2024; Zhu et al., 2024; Park et al., 2023; Qian et al., 2025a).
Yet demonstrating that LLMs can perform economic tasks leaves open the question of how they do so. Imas et al. (2025) study an experimental marketplace in which humans write prompts for AI agents that negotiate on their behalf. They find that, contrary to predictions of homogenized outcomes, agentic interactions exhibit greater dispersion than human-to-human negotiations. This suggests that prompts are not neutral instructions but carriers of human heterogeneity. Our work complements theirs: while they focus on how human characteristics propagate through delegation, we focus on comparing how different agent types behave when facing identical induced preferences, isolating strategic differences rather than principal heterogeneity.
These findings underscore a broader point: AI agents warrant behavioral study in their own right. Rahwan et al. (2019) argue that the behavioral study of AI agents necessitates its own scientific research agenda. This is particularly critical in mixed-motive and multi-agent settings, where alignment is not static but dynamically negotiated. Recent work on virtual bargaining suggests that effective coordination requires agents to mentally simulate implicit social contracts rather than solely maximizing reward functions (Chater, 2023). Furthermore, Bansal et al. (2019) highlight a “performance-compatibility tradeoff”: updates that increase an agent’s theoretical accuracy can paradoxically degrade team performance if they violate the user’s mental model of the agent’s error boundaries.
@@
More generally, a broad theme in recent findings is that appropriate deployment depends not only on system capabilities (Jahani et al., 2025), but also on alignment with user expectations and social norms (Mozannar et al., 2023; Agarwal et al., 2023; Mullainathan and Obermeyer, 2021; Palminteri et al., 2024; Kapania et al., 2022; Qian and Wexler, 2024). Addressing these alignment gaps requires evaluating agents in dynamic contexts with social consequences—a need recognized by recent calls for rigorous methods to steer and assess agent behavior (Raman et al., 2024; Goktas et al., 2025). We contribute to this agenda by studying how humans, LLM-based agents, and Bayesian models behave in a dynamic bargaining task, offering a comparative lens on alignment in strategic social interactions.
3. The Bargaining Game
Figure 1 provides an overview of the game. Below, we describe the setup, information structure, and gameplay.
Setup.
A game consists of 3 players (any $n>2$ suffices) and $k$ distinct types of colored chips, where $k\in\{2,3,4\}$ varies across experimental conditions. [^3] Game complexity is increasing in $k$. Proof provided in Appendix A.1. Players begin with 10 chips of each type. One chip type is always green, which serves as a shared numeraire publicly valued at \$0.50 by all players. The remaining chip types (drawn from red, blue, and purple, depending on $k$) have private valuations: each player independently draws a value for each non-green chip type uniformly from [\$0.10, \$1.00]. While players do not know the valuations other players hold for non-green chips, they do know the distribution from which these values are drawn. Crucially, this setup provides players with an incentive to trade because the possible mutual gains may be substantial. For example, if one player values blue chips at \$0.20 and another values them at \$0.90, trading blue chips for green chips (the shared numeraire) allows both players to capture part of the \$0.70 valuation gap.
@@ 3. The Bargaining Game
Players take turns in a fixed and initially randomized order, cycling through until each player has proposed three times (9 total proposals for 3 players). On their turn, a player proposes a trade: offering chip(s) of one color in exchange for chip(s) of another. The non-proposing players simultaneously decide whether to accept or reject. If both accept, one is chosen at random to complete the trade. Players may only propose or accept trades that they can fulfill with their current inventory. At the end of the game, each player’s payout equals the difference in value between their final and initial chip holdings, floored at zero. Throughout the game, players can see the number of chips each player holds, although not their opponents’ valuations, along with a live history of all previous proposals and trades.
4. A Pareto-Optimal Upper Bound
Next, we provide a theoretical upper bound analysis to benchmark agent performance in our empirical results. Due to the complexity of identifying a tractable, exact equilibrium in this setting, we instead compute a Pareto-optimal allocation via linear programming.
4.1. Notation
Consider a game M = (I, G, v, a) with a set of agents I and chips G. Each agent $i$ has a personal valuation $v_{ig}$ for each good $g$. These valuations are static, independent across agents, and complete for each agent-good pair. All agents have an initial allocation $a^{0}_{ig}$ of each chip, which indicates the amount of chip $g$ allocated to agent $i$. Given these allocations, an agent’s welfare $w_{i}$ is the sum of their valuations weighted by their allocations: $w_{i}=\sum_{g\in G}v_{ig}a_{ig}$. The total welfare w is thus the sum of all individual welfare: $\textbf{w}=\sum_{i\in I}w_{i}=\sum_{i\in I}\sum_{g\in G}v_{ig}a_{ig}$. An individual’s surplus (and corresponding final payout) is defined as the difference between their final and initial welfare.
4.2. Efficient Allocations
In our game, agents trade chips to improve their surplus. Given their initial endowments $a^{0}_{ig}$, mutually beneficial trades may exist that increase total welfare. We provide an upper bound to the obtainable total welfare, assuming the Pareto condition that no player ends with worse utility than they began with.
@@ 4.2. Efficient Allocations
An optimal allocation of chips is any set of allocations $\mathbf{A^{*}}=\{a^{*}_{ig}\ |\ i\in I,g\in G\}$ that solves:
@@
Under a centralized setting with full information, $\mathbf{A^{*}}$ represents a Pareto-efficient allocation that maximizes aggregate utility without making any agent worse off. While this optimal allocation $\mathbf{A^{*}}$ may not be unique, the maximum total welfare $\mathbf{w^{*}}$ is unique. This optimization problem can be solved efficiently using interior point methods. Its computational complexity is bounded by $O\bigl((|I||G|)^{3.5}\bigr)$. Note that this upper bound is unlikely to be achieved in practice, as no mechanism for decentralized bargaining with private information can guarantee both Pareto efficiency and incentive compatibility (Myerson and Satterthwaite, 1983). [^4] See Appendix A.2 for a proof that players have no dominant strategy in our game.
5. Experimental Setup
We implemented the game using Deliberate Lab (Qian et al., 2025b), an open-source platform for conducting multi-agent experiments. Our setup ensures full reproducibility, allowing for exact replication of game mechanics for benchmarking of future LLMs. Figures 6 and  B.1 provide visualizations of the game interface on Deliberate Lab.
In our experimental design, human participants were randomly assigned to groups of three to complete either the 2-chip, 3-chip, or 4-chip game. Humans also completed an end-of-task survey that asked them qualitative questions about their experience playing the game. [^5] Full survey questions are provided in Appendix C.
These human game instances—defined by the unique initial chip endowments within a group—were then replicated in isolation using LLM and Bayesian agents. To facilitate a direct performance comparison, all groups remained homogeneous; no interactions occurred between different agent types (e.g., humans played exclusively against humans, and LLMs against LLMs).
5.1. Human Subjects